OpenStudy (anonymous):

On a blueprint, the scale indicates that 7 cm represent 16 ft. What is the length of a room that is 9.8 cm long and 5 cm wide on the blueprint?

OpenStudy (nubeer):

1 cm = 16/7 ft so length = 9.8 * (16/7) = ? width = 5*(16/7)= ?

OpenStudy (nincompoop):

Do you understand nubeer's solution?

OpenStudy (anonymous):

22.4 11.4 @nubeer

OpenStudy (anonymous):

22.4 11.4 @nubeer

OpenStudy (anonymous):

22.4 11.4 @nubeer

OpenStudy (nincompoop):

This means that if: 16 ft = 7 cm We need to divide 16 by 7 to figure out how many ft there are in a 7 cm. And since the dimensions of the room to scale is 9.8 by 5 in cm, we can multiply the quotient by the length and width respectively.

OpenStudy (anonymous):

ok so are my answers correct ?

OpenStudy (nincompoop):

No they are not correct because you did not answer them by your self. The calculation were precise though

OpenStudy (nincompoop):

I don't even know if you understand what you did. You don't units in your final answer...

OpenStudy (anonymous):

its just a pretest... to other things.. thank you for your help and your smart mouth !